THE ZEEMAN EFFECT IN THE OPTICAL SPECTRUM OF MANGANESE MONOHYDRIDE: MnH^{a}.

\underline{JAMIE GENGLER} and TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ, USA 85287; MICHAEL STOLL, Molecular Physics, Fritz-Haber-Institut der Max-Planck-Gesselschaft, Faradayweg 4-6, 14195 Berlin, Germany.

The only successful demonstration of buffer-gas cooling and magnetic trapping of paramagnetic molecules was that performed by Doyle and co-workers in 1998^{b}. It is anticipated that the relatively large rotational spacing and small spin-spin and spin-rotation interaction in the X^{7}Σ^{+} (v = 0) state of MnH is a signature of a favorable ratio of elastic to inelastic collisions with helium atoms which is required for buffer-gas cooling and trapping. Furthermore, the large magnetic moment of the N = 0, X^{7}Σ^{+} (v = 0) state of MnH enables one to create a deep trap, and this allows a large enough trapping time to be able to isolate the sample from its environment by cryopumping the helium away. Here we report on the magnetic properties of MnH studied as a necessary preliminary to our use of it in buffer-gas cooling and magnetic trapping experiments. The Zeeman effect in the P_{1}(0) branch feature (ν = 17568.3536 cm^{−1}) of the A^{7}Π−X^{7}Σ^{+} (0,0) band system of MnH was selected for Zeeman measurements. The field free spectra is complicated but has been thoroughly analyzed^{c}’^{d}’^{e}.

aSupported by the National Science Foundation-Exp. Phys. Chem.

bJ.D. Weinstein, R. deCarvalho, T. Guillet, B. Friedrich, and J.M. Doyle, Nature 395, 148 (1998).

cT.D. Varberg, R.W. Field, and A.J. Merer, J. Chem. Phys. 95(3), 1563-1576 (1991).

dT.D. Varberg, J.A. Gray, R.W. Field, and A.J. Merer, J. Mol. Spectrosc. 156, 296-318 (1992).

eJ. Gengler, T.C. Steimle, J. Harrison, and J.M. Brown, J. Mol. Spectrosc. 241, 192 (2007).